OpenStudy (anonymous):

When –3 is multiplied by the sum of a number and 8, the result is –21. Find the number. A. 3 B. –1 C. 8 D. –2

OpenStudy (anonymous):

go for it @ryan123345

OpenStudy (ryan123345):

haha

OpenStudy (ryan123345):

-3(x+8)=-21

OpenStudy (ryan123345):

-3x -24=-21 -3x=3 x=-1